OpenStudy (anonymous):

6/5w+2/5w=-1/5 - 1/5

OpenStudy (anonymous):

w=?

OpenStudy (aroub):

8/5w=-2/5 40=-10w w=-4

OpenStudy (anonymous):

6/5w + 2/5w = -1/5 -1/5 6/5w + 2/5w = (6+2)/5w = 8/5w -1/5 -1/5 = (-1-1)/5 = -2/5 8/5w = -2/5 Cross multiply 8*5 = -2*5w 40 = -10w Divide by -10 -4 = w

OpenStudy (anonymous):

it was -1/4

OpenStudy (anonymous):

I'm guessing you need parenthesis: \[(6/5)w+(2/5)w=-1/5-1/5\] (8/5)w=-2/5 w=-1/4

OpenStudy (aroub):

oohhhh!! ssoorryy... :(, yes i guess mcgyver is right! :)

OpenStudy (anonymous):

The error was the OP. It happens frequently.